OpenStudy (anonymous):

Graph for a medal? Graph the inverse of y=3x+3

zepdrix (zepdrix):

Do you understand what the graph of `this` function looks like? The inverse would be that function reflected over the y=x diagonal. Or we can go through the algebra steps to find the inverse, that might be a little easier.

OpenStudy (anonymous):

sure

zepdrix (zepdrix):

\[\large y=3x+3\]We start by swapping y and x.\[\large x=3y+3\]This new y represents the inverse function, let's solve for our y. We'll subtract 3 from each side,\[\large x-3=3y\]Divide each side by 3,\[\large \frac{x}{3}-1=y\]Which we'll write this way,\[\large y=\frac{1}{3}x-1\] So here is our inverse function, any confusion in those steps?

OpenStudy (anonymous):

none at all

zepdrix (zepdrix):

To properly graph the function all we need to do is plot 2 points then draw a line through them. We can see that this function has a y-intercept of -1. So we'll pass through the point \(\large (0,-1)\). What's another nice easy point we could plot? If we let x=3 it gives us umm.. looks like y=0 right? So there another point we can place on our graph \(\large (3,0)\). Understand how to draw your line with those two points? :)

OpenStudy (anonymous):

like this?